# Huairou District, Beijing 101408, P. R. China

### 27 Feb 2024 v1

@@ 1 Introduction
Gravitational waves (GWs) are the latest observationally confirmed prediction of Einstein’s gravity LIGOScientific:2016aoc . Pure Einstein gravity (in the absence of matter fields) admits GW solutions. In 4 dimensional ($4d$) spacetimes, which will be the focus of this work, GWs have a 2 dimensional ($2d$) wavefronts. One can recognize two general classes of GWs, planar wavefront GWs (pp-waves) kundtpp-wave and those with compact wavefronts. The famous solutions of the former are Brinkmann waves Brinkmann:1925fr and the latter are the Robinson-Trautman (RT) Robinson:1960zzb ; Robinson:1962zz solutions. See a recent paper by Roger Penrose Penrose:2024fdd for some historical remarks on these solutions. In this work, we explore further the class of RT solutions.
The RT solutions are solutions to pure Einstein gravity admitting a shear and twist-free congruence of null geodesics with non-zero expansion Stephani:2003tm . [^1] In contrast, pp-waves admit shear and twist-free null geodesic congruence with zero expansion kundt1962study . The Goldberg-Sachs theorem Goldberg-Sachs states that any vacuum Einstein solution with shear-free null congruence is algebraically special and falls into Petrov type II, III, N, D, or O classes. Einstein field equations for the RT family reduce to a single parabolic fourth-order differential equation, the RT equation. RT equation, as pointed out by K. Tod Tod_1989 , is the Calabi flow equation calabi1982extremal ; calabi2002space . The RT solutions have been constructed in de Sitter (dS) or anti-de Sitter (AdS), as well as flat backgrounds. From a physical perspective, the RT metrics can be interpreted as representing a system that is isolated and emits gravitational radiation which typically settles down to a Schwarzschild-like black hole. In this regard, this class of solutions has a specific property that allows one to study black holes and gravitational waves simultaneously.
@@
In a parallel line of developments, starting with the seminal work of Bondi et al and Sachs Bondi:1962px ; Sachs:1962 , asymptotic symmetries of spacetimes which may involve GWs have been explored. The variation in the BMS charges (in asymptotically flat spacetimes) Barnich:2009se ; Strominger:2017zoo has been linked to the gravitational memory effects Strominger:2014pwa ; Pasterski:2015tva . In these analyses, certain falloff behavior for the metric of spacetime (and hence for GWs) are considered and moreover, the topology of the spatial part of the null infinity is assumed to be a two-sphere. The class of RT solutions do not necessarily respect either of these two assumptions. In addition, we may have RT solutions in asymptotic (A)dS backgrounds. Extension of asymptotic symmetry analysis beyond standard BMS ones has been also considered, see e.g. Compere:2020lrt ; Geiller:2022vto and Fernandez-Alvarez:2020hsv ; Fiorucci:2020xto ; Poole:2021avh ; Perez:2022jpr ; Bonga:2023eml ; Geiller:2024amx ; Mao:2018xcw ; Mao:2019ahc . The boundary symmetry analysis is not limited to asymptotic regions of spacetime and has been extended to any boundary, any codimension 1 surface. In particular, boundary symmetries for null or timelike boundaries have been analyzed Adami:2020amw ; Adami:2021nnf ; Adami:2021kvx ; Adami:2023wbe ; Sheikh-Jabbari:2022mqi ; Chandrasekaran:2018aop ; Chandrasekaran:2020wwn ; Ciambelli:2023mir ; Donnelly:2020xgu .
In this paper, our focus is on exploring the solution space of $4d$ RT metrics in the presence of a generic causal boundary. We construct the solution space and analyze its inherent symmetries, the symplectic structure, and the associated surface charges. This investigation is a continuation of our previous studies, where we analyzed symmetries and charges near a generic null surface for $2d$ or $3d$ gravity theories, as discussed in Adami:2020ugu . This analysis was later extended to $D$ dimensional Einstein gravity Adami:2021nnf . In Adami:2021kvx , we established that the boundary degrees of freedom, associated with a generic codimension one null surface in $D$ dimensional pure Einstein gravity, naturally lend themselves to a thermodynamic description. Furthermore, we considered the causal (timelike or null) boundaries for $2d$ and $3d$ gravity theories in Adami:2022ktn , and presented its hydrodynamic description in Adami:2023fbm .
@@
We extend the RT solution to accommodate the presence of a codimension 1 boundary which we typically choose it to be a causal hypersurface in a background with generic cosmological constant $\Lambda$ and allow for the wavefront to be a smooth compact $2d$ surface of arbitrary topology. Our solution space beside the usual time dependent mass and RT scalar mode which describes the time dependence and shape of the RT wavefront and the genus of the wavefront is described by two scalar boundary (soft) modes and a vector field that indicates the local angular velocity of causal hypersurfaces. The two boundary scalar modes are described by two generic functions on the codimension 1 boundary and the vector field which is defined on a codimension 1 surface, satisfies the conformal Killing equations (CKE) over the wavefront.
@@
Next, we study the on-shell symplectic form over the solution phase space and then the surface charges associated with the boundary symmetries. The symplectic form consists of 2 parts: “thermodynamic”, “codimension 2” (boundary symplectic form). The bulk RT mode do not appear in the symplectic form, as discussed in Adami:2023wbe . We find that only particular combinations, contingent upon the chosen slicing Adami:2020ugu ; Adami:2021nnf ; Adami:2022ktn , result in non-zero charges for the given boundary symmetries. We demonstrate the existence of a Heisenberg slicing, wherein the surface charge algebra adopts the structure of a Heisenberg algebra, with the time-independent (ADM) mass serving as a Casimir of the algebra. Furthermore, we study charges associated with superrotation symmetry generators and discuss that there is no independent charge (other than the charges appearing in the Heisenberg algebra) associated with them.
Lastly, we study the generalized RT solution within the fluid/gravity correspondence framework. Considering the on-shell RT symplectic form and surface charges, we construct the RT energy-momentum tensor for both asymptotically AdS and flat spacetimes. In the context of the AdS case, our findings extend the analysis presented in BernardideFreitas:2014eoi ; Ciambelli:2017wou . For the flat case, we construct a null energy-momentum tensor, utilizing the Carrollian structure of the null infinity. We then consider its conservation through a null (Carrollian) version of the Levi-Civita connection. The conservation of this energy-momentum tensor results in the RT equation. Interestingly, we demonstrate that the RT equation can be written in terms of a current conservation equation, in addition to its tensor representation. This is applicable for both AdS and flat cases.
@@ 2 Constructing solution space
where $v$ is advanced time coordinate, $r$ is the radial coordinate and $x^{A}\,,A=1,2,$ are coordinates on a compact two dimensional surface $\Sigma_{g}$. We fix the coordinate such that $\eta$ does not depend on radial coordinate while $V$, $U^{A}$ are generic functions on the spacetime. We study class of geometries with the following properties:
@@ 2.1 Various wavefront topologies
In this case we deal with a hyperboloid $\mathbb{H}^{2}$. $\mathbb{H}^{2}$ has uniform negative curvature, is non-compact and has isometry $so(2,1)$. One makes it a compact space by orbifolding it by freely-acting discrete subgroups of its isometry group. Depending on the choice of the subgroup we obtain a smooth compact manifold with genus $\mathrm{g}>1$. The action of this orbifolding can be reflected in the range of coordinates parameterizing the $H^{2}$, $z=\tanh{(\psi/2)}\,e^{i\phi}$, where $0\leq\psi<\cosh^{-1}{(2\mathrm{g}-1)}$ and $\phi\in[0,2\pi]$. The area element is
@@ 2.3 More on solutions to field equations
where ${\cal D}_{\beta}:=\partial_{\beta}-\mathcal{L}_{\tilde{\mathcal{U}}}$, and $\tilde{\lambda}_{b}:=\tilde{\lambda}_{b}(z,\bar{z})$ are integration constants which may be viewed as $\tilde{\lambda}_{b}(z,\bar{z})=\tilde{\lambda}(v=v_{b};z,\bar{z})$. That is, (2.24) specifies $\tilde{\lambda}(v;z,\bar{z})$ up to a codimension 2 function. RT equation (2.30) is compatible with various wavefront topologies. So, along with the shape function $\tilde{\lambda}_{b}(z,\bar{z})$ one should also specify the topology of the wavefront and in particular its genus $\mathrm{g}$.
To summarize, the solution space is specified through codimension 1 functions $\Omega(v;z,\bar{z}),\mathrm{P}(v;z,\bar{z})$; codimension 2 functions ${\cal U}(v;z),{\bar{\cal U}}(v;{\bar{z}})$, $\tilde{\lambda}_{b}(z,\bar{z})$; ${\cal R}$ as function of only $v$ and finally the constant ${\cal M}_{\circ}$ and the genus of wavefront $\mathrm{g}$. We will address the question that which of these functions are physical and which are pure gauge and can be removed upon proper diffeomorphisms in section 4 where we analyze the surface/boundary charges.
@@ 3 Symplectic structure, general discussion
give rise to two terms, the bulk, and the boundary terms. The bulk term gives us equations of motion $G_{\mu\nu}+\Lambda g_{\mu\nu}=0$ and we refer to the boundary term as symplectic potential Lee:1990nz ; Wald:1999wa
@@
The thermodynamical part is the one without time and space integrals. The boundary term $\boldsymbol{\Omega}^{\mathrm{b}}_{\text{\tiny reg.}}$, as in Adami:2023wbe , involves a codimension 2 spatial integral and $\boldsymbol{\Omega}^{\mathrm{s}}_{\text{\tiny reg.}}$ vanishes unless we allow for “string-like” sources. In this work, we focus on cases with no such sources and leave the discussion on string-like sources to future work. Compared to Adami:2023wbe , our symplectic form does not involve a “bulk” term involving a codimension 1 spatial integral. The reason is in our case we have turned off the “tensor modes” (gravitational waves with non-compact wavefronts), $N_{AB}$ modes in the notation of Adami:2023wbe , in our solution space. As discussed there, the RT modes, correspond to the kernel of the Carrollian solution space and hence drop out of the symplectic form once we turn off $N_{AB}$.
@@ 6.1 Geometric structure
We shall now define the Levi-Civita-Carroll connection. To this end, let us consider the metric-connection compatibility condition,
ˇ$\check{\nabla}_{c}\gamma_{ab}=\partial_{c}\gamma_{ab}-\check{\Gamma}^{d}_{bc}\gamma_{ad}-\check{\Gamma}^{d}_{ac}\gamma_{bd}=0\qquad\Rightarrow\qquad\gamma_{cd}\check{\Gamma}^{d}_{ab}=\frac{1}{2}\left(\partial_{a}\gamma_{bc}+\partial_{b}\gamma_{ac}-\partial_{c}\gamma_{ab}\right)\,.$ (6.6)
Since $\gamma_{ab}$ is not invertible, we shall contract both sides of the equation above with $h^{ec}$ instead. This yields Levi-Civita-Carroll connections
$\breve{\Gamma}^{c}_{ab}:=P^{c}{}_{d}\check{\Gamma}^{d}_{ab}=\frac{1}{2}h^{cd}\left(\partial_{a}\gamma_{bd}+\partial_{b}\gamma_{ad}-\partial_{d}\gamma_{ab}\right)\,,$ (6.7)
@@
We note that ˘$\breve{\nabla}_{c}h^{ab}=0$. While we use the prescription of Bergshoeff:2017btm to compute the Riemann tensor, the connection defined above is a different one. The connection we defined here reduces to spatial and temporal connections defined in Bagchi:2023rwd when we set $\mathcal{U}^{A}=0$.
@@
Given the Levi-Civita-Carroll connection (6.7), we can define the corresponding null Carroll-Riemann tensor
@@ 7 Discussion
We have constructed the most general family of solutions in vacuum Einstein-$\Lambda$ theory which admits a twist and shear-free null geodesic congruence with non-zero expansion. That is, we have extended the Robinson-Trautman (RT) solution to incorporate not only the ADM mass and RT scalar mode $\lambda(v,z,\bar{z})$ (which parameterizes the time-dependent shape of the compact wavefront) but also two boundary modes and a vector field indicative of the local angular velocity of causal hypersurfaces. The analysis has revealed that the vector field, defined on a codimension one surface, satisfies the conformal Killing equations on the wavefront. Within this class of solutions, the wavefront, defined as a smooth compact $2d$ surface, exhibits versatility, allowing for an arbitrary genus $\mathrm{g}$. Additionally, the cosmological constant $\Lambda$ can vary, being positive, zero, or negative.
@@
Moreover, we have thoroughly explored the inclusion of conventional superrotations within the spectrum of surface charges. We have discussed that assuming the smoothness of the wavefront $\Sigma$ only there are 6 charges forming $so(3,1)$ algebra in this sector and moreover, these 6 charges are not independent of the Heisenberg boundary charges.

### 14 Mar 2024 v2

@@ 1 Introduction
Gravitational waves (GWs) are the latest observationally confirmed prediction of Einstein’s gravity LIGOScientific:2016aoc . Pure Einstein gravity (in the absence of matter fields) admits GW solutions. In 4 dimensional ($4d$) spacetimes, which will be the focus of this work, GWs have a 2 dimensional ($2d$) wavefronts. One can recognize two general classes of GWs, planar wavefront GWs (pp-waves) kundtpp-wave and those with compact wavefronts. The famous solutions of the former are Brinkmann waves Brinkmann:1925fr and the latter are the Robinson-Trautman (RT) Robinson:1960zzb ; Robinson:1962zz solutions. See a recent paper by Roger Penrose Penrose:2024fdd for some historical remarks on these solutions. In this work we explore further the class of RT solutions.
The RT solutions are solutions to pure Einstein gravity admitting a shear and twist-free congruence of null geodesics with non-zero expansion Stephani:2003tm . [^1] In contrast, pp-waves admit shear and twist-free null geodesic congruence with zero expansion kundt1962study . The Goldberg-Sachs theorem Goldberg-Sachs states that any vacuum Einstein solution with shear-free null congruence is algebraically special and falls into Petrov type II, III, N, D or O classes. Einstein field equations for the RT family reduce to a single parabolic fourth-order differential equation, the RT equation. RT equation, as pointed out by K. Tod Tod_1989 , is the Calabi flow equation calabi1982extremal ; calabi2002space . The RT solutions have been constructed in de Sitter (dS) or anti-de Sitter (AdS), as well as flat backgrounds. From a physical perspective, the RT metrics can be interpreted as representing a system that is isolated and emits gravitational radiation which typically settles down to a Schwarzschild-like black hole. In this regard, this class of solutions has a specific property that allows one to study black holes and gravitational waves simultaneously.
@@
In a parallel line of developments, starting by the seminal work of Bondi et al and Sachs Bondi:1962px ; Sachs:1962 , asymptotic symmetries of spacetimes which may involve GWs have been explored. The variation in the BMS charges (in asymptotically flat spacetimes) Barnich:2009se ; Strominger:2017zoo has been linked to the gravitational memory effects Strominger:2014pwa ; Pasterski:2015tva . In these analysis, certain falloff behavior for the metric of spacetime (and hence for GWs) are considered and moreover, the topology of the spatial part of the null infinity is assumed to be a two-sphere. The class of RT solutions do not necessarily respect either of these two assumptions. In addition, we may have RT solutions in asymptotic (A)dS backgrounds. Extension of asymptotic symmetry analysis beyond standard BMS ones has been also considered, see e.g. Compere:2020lrt ; Geiller:2022vto and Fernandez-Alvarez:2020hsv ; Fiorucci:2020xto ; Poole:2021avh ; Perez:2022jpr ; Bonga:2023eml ; Geiller:2024amx ; Mao:2018xcw ; Mao:2019ahc . The boundary symmetry analysis is not limited to asymptotic regions of spacetime and has been extended to any boundary, any codimension 1 surface. In particular, boundary symmetries for null or timelike boundary have been analyzed Adami:2020amw ; Adami:2021nnf ; Adami:2021kvx ; Adami:2023wbe ; Sheikh-Jabbari:2022mqi ; Chandrasekaran:2018aop ; Chandrasekaran:2020wwn ; Ciambelli:2023mir ; Donnelly:2020xgu .
In this paper, our focus is on exploring the solution space of $4d$ RT metrics in presence of a generic causal boundary. We construct the solution space and analyze its inherent symmetries, the symplectic structure, and the associated surface charges. This investigation is a continuation of our previous studies, where we analyzed symmetries and charges near a generic null surface for $2d$ or $3d$ gravity theories, as discussed in Adami:2020ugu . This analysis was later extended to $D$ dimensional Einstein gravity Adami:2021nnf . In Adami:2021kvx , we established that the boundary degrees of freedom, associated with a generic codimension one null surface in $D$ dimensional pure Einstein gravity, naturally lend themselves to a thermodynamic description. Furthermore, we considered the causal (timelike or null) boundaries for $2d$ and $3d$ gravity theories in Adami:2022ktn , and presented its hydrodynamic description in Adami:2023fbm .
@@
We extend the RT solution to accommodate the presence of a codimension 1 boundary which we typically choose to be a causal hypersurface in a background with generic cosmological constant $\Lambda$ and allow for the wavefront to be a smooth compact $2d$ surface of arbitrary topology. Our solution space beside the usual time dependent mass and RT scalar mode which describes the time dependence and shape of the RT wavefront and the genus of the wavefront is described by two scalar boundary (soft) modes and a vector field that indicates the local angular velocity of causal hypersurfaces. The two boundary scalar modes are described by two generic functions on the codimension 1 boundary and the vector field which is defined on a codimension 1 surface, satisfies the conformal Killing equations (CKE) over the wavefront.
@@
Next, we study the on-shell symplectic form over the solution phase space and then the surface charges associated with the boundary symmetries. The symplectic form consists of 2 parts: “thermodynamic”, “codimension 2” (boundary symplectic form). The bulk RT mode does not appear in the symplectic form, as discussed in Adami:2023wbe . We find that only particular combinations, contingent upon the chosen slicing Adami:2020ugu ; Adami:2021nnf ; Adami:2022ktn , result in non-zero charges for the given boundary symmetries. We demonstrate the existence of a Heisenberg slicing, wherein the surface charge algebra adopts the structure of a Heisenberg algebra, with the time-independent (ADM) mass serving as a Casimir of the algebra. Furthermore, we study charges associated with superrotation symmetry generators and discuss that there is no independent charge (other than the charges appearing in the Heisenberg algebra) associated with them.
Lastly, we study the generalized RT solution within the fluid/gravity correspondence framework 222 For a review of prior efforts in this direction with various methodologies see Ciambelli:2018wre ; Miskovic:2023zfz ; Campoleoni:2023fug ; Ciambelli:2018xat ; Ciambelli:2020ftk ; Campoleoni:2022wmf ; Mittal:2022ywl .. Considering the on-shell RT symplectic form and surface charges, we construct the RT energy-momentum tensor for both asymptotically AdS and flat spacetimes. In the context of the AdS case, our findings extend the analysis presented in BernardideFreitas:2014eoi ; Ciambelli:2017wou . For the flat case, we construct a null energy-momentum tensor, utilizing the Carrollian structure of the null infinity. We then consider its conservation through a Carroll-Affine connection. The conservation of this energy-momentum tensor results in the RT equation. Interestingly, we demonstrate that the RT equation can be written in terms of a current conservation equation, in addition to its tensor representation. This is applicable for both AdS and flat cases.
@@ 2 Constructing solution space
where $v$ is advanced time coordinate, $r$ is the radial coordinate and $x^{A}\,,A=1,2,$ are coordinates on a compact two dimensional surface $\Sigma_{g}$. We fix the coordinate such that $\eta$ does not depend on radial coordinate while $V$, $U^{A}$ are generic functions on the spacetime. We study a class of geometries with the following properties:
@@ 2.1 Various wavefront topologies
In this case we deal with a hyperboloid $\mathbb{H}^{2}$. $\mathbb{H}^{2}$ has uniform negative curvature, is non-compact and has isometry $so(2,1)$. One makes it a compact space through orbifolding it by freely-acting discrete subgroups of its isometry group. Depending on the choice of the subgroup we obtain a smooth compact manifold with genus $\mathrm{g}>1$. The action of this orbifolding can be reflected in the range of coordinates parameterizing the $H^{2}$, $z=\tanh{(\psi/2)}\,e^{i\phi}$, where $0\leq\psi<\cosh^{-1}{(2\mathrm{g}-1)}$ and $\phi\in[0,2\pi]$. The area element is
@@ 2.3 More on solutions to field equations
where ${\cal D}_{\beta}:=\partial_{\beta}-\mathcal{L}_{\tilde{\mathcal{U}}}$, and $\tilde{\lambda}_{b}:=\tilde{\lambda}_{b}(z,\bar{z})$ are integration constants which may be viewed as $\tilde{\lambda}_{b}(z,\bar{z})=\tilde{\lambda}(v=v_{b};z,\bar{z})$. That is, (2.24) specifies $\tilde{\lambda}(v;z,\bar{z})$ up to a codimension 2 function. RT equation (2.30) is compatible with various wavefront topologies. So, along with the shape function $\tilde{\lambda}_{b}(z,\bar{z})$ one should also specify topology of the wavefront and in particular its genus $\mathrm{g}$.
To summarize, the solution space is specified through codimension 1 functions $\Omega(v;z,\bar{z}),\eta(v;z,\bar{z})$; codimension 2 functions ${\cal U}(v;z),{\bar{\cal U}}(v;{\bar{z}})$, $\tilde{\lambda}_{b}(z,\bar{z})$; ${\cal R}$ as function of only $v$ and finally the constant ${\cal M}_{\circ}$ and the genus of wavefront $\mathrm{g}$. We will address the question that which of these functions are physical and which are pure gauge and can be removed upon proper diffeomorphisms in section 4 where we analyze the surface/boundary charges.
@@ 3 Symplectic structure, general discussion
give rise to two terms, the bulk and the boundary terms. The bulk term gives us equations of motion $G_{\mu\nu}+\Lambda g_{\mu\nu}=0$ and we refer to the boundary term as symplectic potential Lee:1990nz ; Wald:1999wa
@@
The thermodynamical part is the one without time and space integrals. The boundary term $\boldsymbol{\Omega}^{\mathrm{b}}_{\text{\tiny reg.}}$, as in Adami:2023wbe , involves a codimension 2 integral and $\boldsymbol{\Omega}^{\mathrm{s}}_{\text{\tiny reg.}}$ vanishes unless we allow for “string-like” sources. In this work, we focus on cases with no such sources and leave the discussion on string-like sources to future work. Compared to Adami:2023wbe , our symplectic form does not involve a “bulk” term involving a codimension 1 spatial integral. The reason is in our case we have turned off the “tensor modes” (gravitational waves with non-compact wavefronts), $N_{AB}$ modes in the notation of Adami:2023wbe , in our solution space. As discussed there, the RT modes, correspond to the kernel of the Carrollian solution space and hence drop out of the symplectic form once we turn off $N_{AB}$.
@@ 6.1 Geometric structure
We define the Carroll-Affine connection as
$\breve{\Gamma}^{c}_{ab}:=\frac{1}{2}h^{cd}\left(\partial_{a}\gamma_{bd}+\partial_{b}\gamma_{ad}-\partial_{d}\gamma_{ab}\right)\,,$ (6.6)
@@
While we use the prescription of Bergshoeff:2017btm to compute the Riemann tensor, the connection defined above differs. The connection we defined here reduces to spatial and temporal connections defined in Bagchi:2023rwd when we set $\varpi=1$ and $\mathcal{U}^{A}=0$. An important point is that connection (6.6) is not compatible with degenerate metric $\gamma_{ab}$, i.e. we have
˘$\breve{\nabla}_{c}\gamma_{ab}=\theta_{l}\,\gamma_{c(a}n_{b)}\,,$ (6.8)
where $\theta_{l}$ is expansion of kernel vector $l^{a}$, that is
$\theta_{l}:=P^{a}{}_{b}\breve{\nabla}_{a}l^{b}=2\,\frac{\mathcal{D}_{v}(\varpi\sqrt{q})}{\alpha\varpi\sqrt{q}}\,.$ (6.9)
In the context of the discussion in section 6.2, it is pertinent to note that while there exists a conformal frame $\varpi=1$ in which the metric and connection are compatible, to express the RT equation as a conservation law for a specified energy-momentum tensor, and to express the symplectic form in terms of the given tensor, we shall fix conformal factor $\varpi$ in particular way. We also note that ˘$\breve{\nabla}_{c}h^{ab}=0$.
@@
Given the Carroll-Affine connection (6.6), we can define the corresponding null Carroll-Riemann tensor
@@ 7 Discussion
We have constructed the most general family of solutions in vacuum Einstein-$\Lambda$ theory which admit a twist and shear-free null geodesic congruence with non-zero expansion. That is, we have extended the Robinson-Trautman (RT) solution to incorporate not only the ADM mass and RT scalar mode $\lambda(v,z,\bar{z})$ (which parameterizes the time-dependent shape of the compact wavefront) but also two boundary modes and a vector field indicative of the local angular velocity of causal hypersurfaces. The analysis has revealed that the vector field, defined on a codimension one surface, satisfies the conformal Killing equations on the wavefront. Within this class of solutions, the wavefront, defined as a smooth compact $2d$ surface, exhibits versatility, allowing for an arbitrary genus $\mathrm{g}$. Additionally, the cosmological constant $\Lambda$ can vary, being positive, zero, or negative.
@@
Moreover, we have thoroughly explored the inclusion of conventional superrotations within the spectrum of surface charges. We have discussed that assuming smoothness of the wavefront $\Sigma$ only there are 6 charges forming $so(3,1)$ algebra in this sector and moreover, these 6 charges are not independent of the Heisenberg boundary charges.